OpenStudy (anonymous):

Classify the following system of equations: 9x − 3y = 12 2y = 6x − 8 a. Consistent and independent b. Inconsistent c. Consistent and dependent

OpenStudy (anonymous):

its consistent with infinite solution

OpenStudy (anonymous):

@matricked : so would that be A or C ?

OpenStudy (anonymous):

i am having the doubt for dependent and independent sols still i suppose is it c

OpenStudy (anonymous):

9x - 3y = 12 6x - 2y = 8 18x - 6y = 24 18x - 6y = 24 These are parallel lines with infinitely many solutions...

OpenStudy (anonymous):

@waterineyes : so would you agree with matricked that its c ?

OpenStudy (anonymous):

it is definitely C

OpenStudy (anonymous):

I think A will be the answer.. But I am not too sure about Dependent and Independent..